If 77.60 g of nitrogen gas are produced in a reaction, what is the volume if the temperature is 48°C and the pressure is 99.5 kPa?
1
Expert's answer
2020-07-14T15:19:00-0400

According to the universal gas law, pV = nRT => V = nRT/p.

n (N2) = 77.60g / 28.01g/mol = 2.77 mol

48oC = 321K

V = (2.77 mol x 8.314J/mol*K x 321K) / 99.5kPa = 74.3 L

Answer: 74.3 L